# Are Multimodal Embeddings Truly Beneficial for Recommendation? A Deep Dive into Whole vs. Individual Modalities

### v2

@@ 5.2.2. The Ineffectiveness of Vision:
6. Discussion
In this section, we focus on models that exhibit distinct behaviors compared with most models, providing a closer examination of the results from RQ1 and RQ2.
6.1. Visually-Sensitive Models
In the RQ2 analysis, we observed that most models make limited use of visual-modality information. Nevertheless, we also identified a subset of models from Group 2 that are highly sensitive to the image modality—when their image embeddings were replaced, their performance degraded substantially. As demonstrated in Table 9, we summarize:
Table 3. Performance Comparison under Image Knockout for Vision-Sensitive Models.
@@ 6.1. Visually-Sensitive Models
Table 4. Performance Impact of Modality Knockouts on BM3, SLMRec, and VBPR.
@@
Our results indicate that the three Group 2 models—GRCN, LGMRec, and MMGCN—exhibit distinct behavior compared to other models: they are highly sensitive to visual features. In practice, this means that deploying these models inherently places greater emphasis on visual information, making them particularly suitable for scenarios with abundant and reliable image data. By contrast, models in other groups show weaker dependence on visual signals, which makes them more resilient when visual information is sparse or noisy.
6.2. Case Study: VBPR’s Failure to Utilize Modality Features
@@
Figure 4. Recall@10 under Different Embedding Replacement Strategies.
@@
6.3. More Exploration of Visual Replacement
Beyond modality knockout replacement, another research direction leverages multimodal large language models (MLLMs) as image captioners. In this approach, the model generates textual captions for images, which are then used as the image modality replacement in multimodal recommendation (Zhang et al., 2024b). We adopt the advanced LLaVA [^4] https://huggingface.co/liuhaotian/llava-v1.5-7b to generate captions, embed them with the same Sentence-BERT. As shown in Table 5 on the Baby dataset, performance generally drops slightly compared with using original image information, and the degradation becomes more pronounced in Appendix Table 13, with over 10% declines for most models. These results suggest that MLLM-derived captions convey limited useful information and can even hinder model learning, performing no better—or worse—than using noise vectors.
Table 5. Performance on Baby Dataset using Llava-generated Image Captions.
Model NDCG@10 Recall@10 Precision@10
BM3 0.0277 (+0.7%) 0.0516 (-0.2%) 0.0057 (+0.0%)
DAMRS 0.0311 (-4.6%) 0.0570 (-5.6%) 0.0063 (-4.5%)
DRAGON 0.0319 (-6.7%) 0.0595 (-4.8%) 0.0065 (-4.4%)
DualGNN 0.0262 (-2.2%) 0.0488 (-0.4%) 0.0054 (+0.0%)
FREEDOM 0.0321 (-2.4%) 0.0605 (-1.9%) 0.0066 (-1.5%)
GRCN 0.0284 (-2.7%) 0.0509 (-4.5%) 0.0056 (-5.1%)
LATTICE 0.0289 (-2.0%) 0.0528 (-1.7%) 0.0058 (-1.7%)
LGMRec 0.0308 (-6.9%) 0.0570 (-4.2%) 0.0063 (-3.1%)
MGCN 0.0350 (+6.4%) 0.0624 (+4.9%) 0.0068 (+4.6%)
MMGCN 0.0189 (-3.6%) 0.0362 (-0.8%) 0.0041 (+0.0%)
PGL 0.0312 (+1.6%) 0.0590 (+1.9%) 0.0065 (+1.6%)
SLMRec 0.0273 (-0.7%) 0.0488 (-0.8%) 0.0054 (+0.0%)
SMORE 0.0369 (+2.8%) 0.0665 (+2.8%) 0.0073 (+2.8%)
7. Conclusion and Future Work
This study critically examines whether current multimodal recommendation models truly leverage all available modality information. Our analysis first reveals that most existing multimodal recommendation models primarily rely on multimodal embeddings, particularly those equipped with sophisticated fusion mechanisms. Surprisingly, some classic baselines with simple fusion strategies, such as VBPR and BM3, exhibit unexpectedly stable performance even when their multimodal embeddings are replaced with meaningless vectors. A closer investigation of individual modalities further shows that text plays a dominant role: most models maintain comparable performance even when image embeddings are removed. We also discuss models whose behavior deviates from this general trend.
These findings have important implications for the multimodal recommendation community. We advocate for modality ablation studies to become a standard evaluation protocol, ensuring that reported improvements over baselines are transparent and credible. Moreover, a deeper understanding of how multimodal embeddings influence multimodal recommendation performance would greatly benefit practitioners in effectively deploying these models. For future model design, researchers should move beyond simple fusion and explore integration strategies where modalities actively influence the construction of user–item and item–item graphs during intermediate computations, supported by training techniques that prevent signal overshadowing.

### v3

@@ 5.2 Individual Modalities (RQ2)
6 Discussion
6.1 Visually-Sensitive Models
In the RQ2 analysis, we observed that most models make limited use of visual-modality information. Nevertheless, we also identified a subset of models from Group 2 that are highly sensitive to the image modality—when their image embeddings were replaced, their performance degraded substantially. Based on the experimental results, we summarize:
Table 2: Performance Comparison under Image Knockout for Vision-Sensitive Models.
@@ 6.1 Visually-Sensitive Models
Our results indicate that the three Group 2 models (GRCN, LGMRec, and MMGCN) exhibit distinct behavior compared to other models: they are highly sensitive to visual features. In practice, this means that deploying these models inherently places greater emphasis on visual information, making them particularly suitable for scenarios with abundant and reliable image data. By contrast, models in other groups show weaker dependence on visual signals, which makes them more resilient when visual information is sparse or noisy.
6.2 Impact of Visual Encoder Choice: Replacing ResNet-50 with ViT
To examine whether our conclusions are specific to the ResNet-50 visual features, we replace the image embeddings with ViT-based representations while keeping the datasets, splits, hyperparameters, and textual embeddings unchanged. Our results show that switching to ViT does not yield uniform gains; instead, the effect is highly model-dependent. For several models, performance changes are marginal, suggesting limited sensitivity to the visual encoder. For example, BM3’s NDCG@10 changes by only +0.7% / +0.5% / 1.2-1.2% on Baby/Clothing/Sports, and SMORE remains similarly stable (+1.1% / 0.9-0.9% / 0.2-0.2%). In contrast, a subset of models exhibits strong encoder sensitivity, with either consistent improvements or notable degradations. GRCN provides a clear positive case, improving NDCG@10 by +1.7% / +3.2% / +2.9% and Recall@10 +1.5% / +2.6% / +3.6% on Baby/Clothing/Sports, indicating that richer ViT features can translate into better ranking accuracy when the architecture effectively exploits visual signals. Meanwhile, several models degrade substantially under ViT on specific datasets. DRAGON drops by 11.1-11.1% in NDCG@10 on Baby and 5.2-5.2% on Sports, MMGCN shows a pronounced decline on Clothing (8.8-8.8% NDCG@10 and 9.4-9.4% Recall@10), and LGMRec decreases by 7.8-7.8% on Clothing and 4.1-4.1% on Sports. This mixed pattern suggests that the benefit of stronger visual representations is not only a function of feature quality, but also depends on how each model integrates visual information, which may require encoder-specific normalization or retuning to remain effective.
Overall, Our results reinforce that visual modality utility is architecture and mechanism dependent: some models can consistently convert stronger visual embeddings into gains, while others remain insensitive or even negatively impacted when the visual feature space changes.
6.3 Case Study: VBPR’s Failure to Utilize Modality Features
Table 3: Performance Impact of Modality Knockouts on BM3, SLMRec, and VBPR.
@@ 6.3 Case Study: VBPR’s Failure to Utilize Modality Features
Figure 4: Recall@10 under Different Embedding Replacement Strategies.
@@
7 Conclusion and Future Work
This study critically examines whether current multimodal recommendation models truly leverage all available modality information. Our analysis first reveals that most existing multimodal recommendation models primarily rely on multimodal embeddings, particularly those equipped with sophisticated fusion mechanisms. Surprisingly, some classic baselines with simple fusion strategies, such as VBPR and BM3, exhibit unexpectedly stable performance even when their multimodal embeddings are replaced with meaningless vectors. A closer investigation of individual modalities further shows that text plays a dominant role: most models maintain comparable performance even when image embeddings are removed. We also discuss models whose behavior deviates from this general trend. These findings have important implications for the multimodal recommendation community.